The point (3,2) lies on thee line 11x - 6y = 21 determine if this point is a solution of the system 11x -6y = 21 9x +8y = 43
Answer by Alan3354(69443)

The point (3,2) lies on thee line 11x - 6y = 21 determine if this point is a solution of the system 11x -6y = 21 9x +8y = 43
----------
Sub the values for x & y into 9x +8y = 43 and see if they make the statement true.